1.consider an individual who has CA degree and is considering venturing into spare parts business rather going for a corporate job which pays 50,000 annually. he invests GH 300,000 in the business which has a bank interest rate of 5%

if sales is GH 100,000; cost of goods is GH40,000; depreciation is GH5000; Utility is GH 4000; advertising is GH10,000; miscellaneous is GH5000

i.determine the accounting cost

ii.economic cost

iii.economic profit

iv.what is the implication of this results to manager of business


2.with appropriate examples discuss the significance of elasticity in economics analysis in ghana


3.explain why the between the average cost an average variable cost keeps declining as output increases an implication for the cost of production

4.profit maximation is arguably the main aim of a business critically evaluate this assertion



1
Expert's answer
2021-05-16T17:39:35-0400

1.

i. The accounting cost is: AC = 40,000 - 5,000 + 4,000 + 10,000 + 5,000 = 54,000.

ii. Economic cost is: EC = 54,000 + 50,000 + 300,000*0.05 = 119,000.

iii. Economic profit is: EP = 100,000 - 119,000 = -19,000.

iv. The implication of this results to manager of business is that he shouldn't start the business.


2.Elasticity is an important economic measure, particularly for the sellers of goods or services, because it indicates how much of a good or service buyers consume when the price changes. When a product is elastic, a change in price quickly results in a change in the quantity demanded.


3.Initially, AVC and ATC decrease as output increases. The law states that at some point, the additional cost incurred to produce one more unit is greater than the additional revenue (or returns) received. At that point, the AVC and ATC start to increase.


4.Profit maximization is the main aim of any business and therefore it is also an objective of financial management. Under the assumptions of given tastes and technology, price and output of a given product under perfect competition are determined with the sole objective of maximising profits.